OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 3. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

\[\log_{16}(32)=x+3\]?

OpenStudy (anonymous):

we have a couple of ways to do this perhaps the easiest way is to rewrite in equivalent exponential form as \[16^{x+3}=32\]

OpenStudy (anonymous):

now we can proceed like in the last question, make the bases the same since \(16=2^4\) and \(32=2^5\) you can write \[2^{4(x+3)}=2^5\] which makes \[4(x+3)=5\]

OpenStudy (anonymous):

okay!!:)))

OpenStudy (anonymous):

you know how to solve this one?

OpenStudy (anonymous):

yes thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

How would you solve it? I'm guessing distribute the 4 to get (4x+12)=5, then subtracting 12 form each side = 4x=-7, dividing by 4, giving me x= -7/4??